OpenStudy (dj3strella):

can someone help me 2. Where is the vertex of the parabola? y = x2 + 2x – 3 Above the x-axis Below the x-axis On the x-axis On the y-axis 3. Use the discriminant to determine how many x-intercepts the graph of the equation has. y = – 4x2 + 4x – 1 zero one two three 4. In which direction does the parabola open? y = 3x2 – 5x – 10 down right left up 5. What are the x-intercepts of the quadratic function? y = x2 + x 2 and 1 –2 and 1 2 and –1

OpenStudy (anonymous):

the first coordinate of the vertex of \(y=ax^2+bx+c\) is always \(-\frac{b}{2a}\) for \(y = x^2 + 2x – 3\) you have \(a=1,b=2\) so the first coordinate is \(-\frac{3}{2}\)

OpenStudy (anonymous):

to find the second coordinate of the vertex, replace \(x\) by \(-\frac{3}{2}\) and see what you get it if is a positive number, then it lies above the \(x\) axis, and if it is a negative number then it lies below the \(x\) axis